You are the fore person of a jury charged with determining the amount of damages to pay to a plaintiff who was injured at work due to her employer’s negligence. Because of her injuries, she lost the previous two years’ salaries ($3,000 per month for the first year and $3,200 per month for the most recent year) and is expected to lose $3,500 per month for the next five years after today.

a. What is the amount of the award you would recommend she receive if the interest rate used to determine the award is 9.0%, compounded monthly? Please calculate today’s value even if she won’t get paid until the trial is over.

b. What is the amount of the award you would recommend she receive today if the interest rate used to determine the award is 5.72%, compounded weekly (assume four weeks per month)?

DO NOT USE EXCEL
Answer #1

TO FIND:-

(a) What is the amount of the award you would recommend she receive if the interest rate used to determine the award is 9.0%, compounded monthly? Please calculate today’s value even if she won’t get paid until the trial is over.

ANSWER:-

9% monthly

=(3000/(9%/12))*((1+9%/12)^12-1)*(1+9%/12)^12+3200/(9%/12)*((1+9%/12)^12-1)+3500/(9%/12)*(1-1/(1+9%/12)^60)=249673.74

******

(b) What is the amount of the award you would recommend she receive today if the interest rate used to determine the award is 5.72%, compounded weekly (assume four weeks per month)?

ANSWER:-

5.72% weekly

4 weeks per month

APR=((1+9%/12)^(12/48)-1)*48

8.975% compounded monthly

=(3000/(8.975%/12))*((1+8.975%/12)^12-1)*(1+8.975%/12)^12+3200/(8.975%/12)*((1+8.975%/12)^12-1)+3500/(8.975%/12)*(1-1/(1+8.975%/12)^60)

=249752.78